OpenStudy (anonymous):

Hi there! Could I have some help solving this Algebra problem? Your classmate is starting a new fitness program. He is planning to ride his bicycle 60 minutes every day. He burns 7 calories per minute bicycling at 11 mph and 11.75 calories per minute bicycling at 15 mph. How long should he bicycle at each speed to burn 600 calories per hour? I'm supposed to use a matrix equation to solve it, but I can't figure it out! Any help would be appreciated u w u

jimthompson5910 (jim_thompson5910):

do you have your equations set up?

OpenStudy (anonymous):

Nope :c That's one of my problems... I can't figure out how to set it up.

jimthompson5910 (jim_thompson5910):

how about your variables?

OpenStudy (anonymous):

Would the variables be c=Calories s=Speed?

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

notice how it asks "How long should he bicycle at each speed to burn 600 calories per hour? "

OpenStudy (anonymous):

Hmmm... Is it speed and time?

OpenStudy (anonymous):

Oh, so it would be speed and calories then?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

it's actually the two different times for each speed so you can say something like x = time devoted to biking at 11 mph y = time devoted to biking at 15 mph

jimthompson5910 (jim_thompson5910):

both times are in minutes

jimthompson5910 (jim_thompson5910):

because ultimately we have to answer the question: How long should he bicycle at each speed to burn 600 calories per hour?

jimthompson5910 (jim_thompson5910):

the answers will be the values of x and y (actual numbers, not just variables)

OpenStudy (anonymous):

Ok, great :) So how would I go about solving this?

jimthompson5910 (jim_thompson5910):

before we solve, we need to set up the equations to solve. It's like setting up the question that we have to answer

jimthompson5910 (jim_thompson5910):

it says "He is planning to ride his bicycle 60 minutes every day." so the first equation is x+y = 60

jimthompson5910 (jim_thompson5910):

it also adds "He burns 7 calories per minute bicycling at 11 mph and 11.75 calories per minute bicycling at 15 mph" because he wants to burn 600 calories, this means the second equation is 7x + 11.75y = 600

OpenStudy (anonymous):

Hold on, before you go any further, I'm supposed to be using matrices to solve this. Shouldn't I put the x and y values into a variable matrix first?

jimthompson5910 (jim_thompson5910):

we'll get there

OpenStudy (anonymous):

Wait nvm, I see where you're going with this.

OpenStudy (anonymous):

Sorry :P

jimthompson5910 (jim_thompson5910):

its ok

jimthompson5910 (jim_thompson5910):

we can multiply both sides of 7x + 11.75y = 600 by 100 to get 7x + 11.75y = 600 100*(7x + 11.75y) = 100*600 700x + 1175y = 6000

jimthompson5910 (jim_thompson5910):

so we have this system of equations x+y = 60 700x + 1175y = 6000 how do we turn this into a matrix equation?

jimthompson5910 (jim_thompson5910):

oh sorry, typo

jimthompson5910 (jim_thompson5910):

700x + 1175y = 6000 should be 700x + 1175y = 60000 forgot a 0

jimthompson5910 (jim_thompson5910):

we really have x+y = 60 700x + 1175y = 60000

OpenStudy (anonymous):

Ok, I've got it written down, and I've set up the matrices. What next?

jimthompson5910 (jim_thompson5910):

what matrix did you get for the left side?

OpenStudy (anonymous):

The coefficient matrix?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

Oh, I got [700 1175] I feel like I'm missing a row...

jimthompson5910 (jim_thompson5910):

yes row 1

OpenStudy (anonymous):

What should I have entered for row one then?

jimthompson5910 (jim_thompson5910):

1 and 1 (the coefficients for x+y)

OpenStudy (anonymous):

Ok, that's what I thought :)

jimthompson5910 (jim_thompson5910):

your 2x2 matrix should be this